OpenStudy (anonymous):

5^{-8x}=16^{x+8} Write the exact answer using base-10 logarithms.

OpenStudy (anonymous):

\[(-8x)(\log5) = (x+8)(\log16)\]

OpenStudy (anonymous):

\[\frac{ -8x }{ x+8 }=\frac{ \log16 }{ \log 5 }\]

OpenStudy (anonymous):

So I'm Stuck there

OpenStudy (tkhunny):

Yes, you are. Don't do that. \((-8x)(log(5)) = (x+8)(log(16)) = xlog(16) + 8log(16)\) Then \((-8x)log(5) - xlog(16) = 8log(16)\) Now what?

OpenStudy (anonymous):

Still stuck. Need to solve for x. I cant see what I could move over from there

OpenStudy (tkhunny):

Have you considered the distributive property? \(x[(-8)log(5) - log(16)] = 8log(16)\)

OpenStudy (anonymous):

I had not. Thank you.

OpenStudy (tkhunny):

Actually, you could have gotten there the way your had it. \(\dfrac{-8x}{x+8} = -8 - \dfrac{64}{x+8} = \dfrac{log(16)}{log(5)}\) It's no less annoying that the other way we did it already. It's just another way.

OpenStudy (anonymous):

Thanks mate

OpenStudy (tkhunny):

Whoops! "+64", not "-64"!